### v1

@@ 3.1.3 Total error
$\quad+\left(\frac{1}{2(k\,k^{\eta})^{3/2}}\cdot\left\|A\right\|_{L^{p,1}(\mathbb{R}^{2})}+\frac{1}{k^{\eta}}\right)\cdot\frac{1}{\varepsilon}\max(C_{\mathbb{R},\Omega},C_{p})\left\|\sum_{j=1}^{N}\eta_{j}\right\|_{W^{1,1}(\mathbb{R}^{2})}$
(16) $\leq C\cdot\left(\left\|\mathcal{A}f\right\|_{L^{p,1}(\mathbb{R}^{2}\backslash\Omega)}+\left\|\mathcal{A}f(0,\cdot)\right\|_{L^{1}((\mathbb{R}^{2}\backslash\Omega)\cap(\left\{0\right\}\times\mathbb{R}))}+\frac{1}{\varepsilon}\left\|\sum_{j=1}^{N}\eta_{j}\right\|_{W^{1,1}(\mathbb{R}^{2})}\right).$
@@
Now, consider a fixed $\varepsilon$. If we only use a single window function, i.e., $N=1$, then the set $\Omega$ can be significantly smaller compared to the case with multiple windows; see Figures 1 and 2 for illustrations. This again sheds lights on the fact that we achieve better stability by using more windows. In this way, we can significantly reduce the magnitude of the first two terms in (16) while keeping the last term constant.
@@ 3.2 Stability of Algorithm 2
In Algorithm 2, the reconstruction is done separately on each $\Omega_{j}$, using $F_{j}$, $F_{j}^{\eta}$ for computing $A$, $A^{\eta}$, respectively. Again, $A$ and $A^{\eta}$ will be set to zero outside of $\Omega$ and we obtain the following error bound for the data noise propagation:
@@
Let $A,A^{\eta}$ be the reconstructions obtained from $F_{j},$ $F_{j}^{\eta},$ respectively, $j=1,\dots,N$, and for some $\varepsilon>0$, let $\Omega_{j}$ be chosen such that $\left|\mathcal{A}h_{n_{j}}(x,y)\right|>\varepsilon$ for all $(x,y)\in\Omega_{j}$ for all $j\in\{1,\dots,N\}$. Then,
$\left\|A-A^{\eta}\right\|_{L^{p,1}(\mathbb{R}^{2})}$$\leq$ $\frac{C_{p}}{\varepsilon}\max_{j=1,\dots,N}\left\|\eta_{j}\right\|_{W^{1,1}(\mathbb{R}^{2})},$
@@
$\left\|A-A^{\eta}\right\|_{L^{p,1}(\mathbb{R}^{2})}\leq\sum_{j=1}^{N}\left\|\mathds{1}_{\Omega_{j}}\frac{T\mathcal{F}(F_{j}-F_{j}^{\eta})}{\mathcal{A}h_{n_{j}}}\right\|_{L^{p,1}(\mathbb{R}^{2})}\leq\sum_{j=1}^{N}C_{j,p}\cdot\left\|\mathds{1}_{\Omega_{j}}\omega\frac{T\mathcal{F}(F_{j}-F_{j}^{\eta})}{\mathcal{A}h_{n_{j}}}\right\|_{L^{\infty}(\mathbb{R}^{2})},$
where $C_{j,p}:=\|\mathds{1}_{\Omega_{j}}/\omega\|_{L^{p,1}(\mathbb{R}^{2})}<\infty$ and instead of (9), we now have the similar bound
$\left\|A-A^{\eta}\right\|_{L^{p,1}(\mathbb{R}^{2})}$ $\leq\sum_{j=1}^{N}\frac{C_{j,p}}{\inf_{\Omega_{j}}|\mathcal{A}h_{n_{j}}|}\cdot\left\|\widetilde{\omega}\,\mathcal{F}\eta_{j}\right\|_{L^{\infty}(\mathbb{R}^{2})}$
$\leq\frac{1}{\varepsilon}\sum_{j=1}^{N}C_{j,p}\cdot\left\|\mathcal{F}\left(\eta_{j}-\frac{\partial}{\partial x}\eta_{j}+\frac{\partial}{\partial y}\eta_{j}\right)\right\|_{L^{\infty}(\mathbb{R}^{2})}$
$\leq\frac{1}{\varepsilon}\sum_{j=1}^{N}C_{j,p}\cdot\left\|\eta_{j}\right\|_{W^{1,1}(\mathbb{R}^{2})}\leq\frac{C_{p}}{\varepsilon}\max_{j=1,\dots,N}\left\|\eta_{j}\right\|_{W^{1,1}(\mathbb{R}^{2})},$
where we have used $\sum_{j=1}^{N}C_{j,p}=C_{p}$ since $\Omega_{1},\dots,\Omega_{N}$ are pairwise disjoint and the Hausdorff-Young inequality. Similar to (10), we can further derive
@@
With this, the propagation of data noise can be estimated by
$\left\|\widetilde{f}-\widetilde{f\,}^{\eta}\right\|_{L^{p}(\mathbb{R})}$ $\leq\left(\frac{1}{2(k\,k^{\eta})^{3/2}}\cdot\left\|A\right\|_{L^{p,1}(\mathbb{R}^{2})}+\frac{1}{k^{\eta}}\right)\cdot\frac{1}{\varepsilon}\max(C_{\mathbb{R},\Omega},C_{p})\max_{j=1,\dots,N}\left\|\eta_{j}\right\|_{W^{1,1}(\mathbb{R}^{2})}.$
@@
$\quad+\left(\frac{1}{2(k\,k^{\eta})^{3/2}}\cdot\left\|A\right\|_{L^{p,1}(\mathbb{R}^{2})}+\frac{1}{k^{\eta}}\right)\cdot\frac{1}{\varepsilon}\max(C_{\mathbb{R},\Omega},C_{p})\max_{j=1,\dots,N}\left\|\eta_{j}\right\|_{W^{1,1}(\mathbb{R}^{2})}$
(17) $\leq C\cdot\left(\left\|\mathcal{A}f\right\|_{L^{p,1}(\mathbb{R}^{2}\backslash\Omega)}+\left\|\mathcal{A}f(0,\cdot)\right\|_{L^{1}((\mathbb{R}^{2}\backslash\Omega)\cap(\left\{0\right\}\times\mathbb{R}))}+\frac{1}{\varepsilon}\max_{j=1,\dots,N}\left\|\eta_{j}\right\|_{W^{1,1}(\mathbb{R}^{2})}\right),$
@@
The third term on the RHS is the only difference between (16) and (17). The former depends on the maximum noise power, while the latter involves the sum of the noise for different window functions. Their differences result from the distinctions between Algorithms 1 and 2 regarding how the multi-window measurements are utilized in the STFT phase retrieval.
@@ 4.1 Discretization
(b) Recovery with a standard Gauss window
Figure 4: (a) Recovery using $40$ fractional Fourier transform of the dilated Gauss windows at various angles; (b) Recovery using a single standard Gauss window. The ground truth is plotted in blue.
@@
Figure 5: (a) Recovery using $6$ Hermite windows of $h_{n}$ where $n=0,10,20,30,40,50$; (b) Recovery using $6$ Hermite windows of $h_{n}$ with the index $n$ randomly selected between 00 and $50$, resulting in a set of indices $\{17,24,29,30,31,49\}$. The ground truth is plotted in blue.
@@ 4.2.1 Chirp signal recovery
As our first inversion example, we consider the chirp signal shown in Figure 3a. As illustrated in Figure 3b, its ambiguity function has extensive support in the time-frequency domain, making it challenging to reconstruct using a single window of the standard Gauss function whose ambiguity function has a very small essential support as seen in Figure 1a. We set $\varepsilon=10^{-3}$ as a cutoff value, as used in Algorithms 1 and 2 to prevent the instability from dividing the phaseless measurements by a small value.
@@
We first use $40$ fractionally transformed dilated Gaussian windows to perform the reconstruction following Algorithm 1, i.e., $\mathcal{F}_{\alpha_{j}}\varphi^{a}$ with $a=15$ and $\alpha_{j}=(j-1)\pi/40$, $j=1,\ldots,40$. The relative error based on formula (20) between the ground truth and the reconstructed signal is $0.0084$. We also use the standard Gauss function $\varphi^{1}$ as a single window function for comparison. The relative error is $0.7928$. After finding the global phase factor, we plot the reconstructions under different measuring window schemes in Figure 4. More measuring windows yield a much more extensive joint coverage of the time-frequency domain, resulting in better signal reconstruction.
Next, we use multiple Hermite windows to perform the recovery. Consider $6$ Hermite function $h_{n}$ with $n\in\{0,10,20,30,40,50\}$. We then follow Algorithm 2 to perform the inversion. The misfit between the reconstructed signal and the ground truth measured by (20) is $0.0025$. On the other hand, we want to test the idea of randomly choosing window functions. We draw $6$ windows uniformly from the zeroth to the fiftieth degree Hermite functions, and use them to generate the phaseless data and perform the STFT phase retrieval following Algorithm 2. We repeat this procedure $100$ times to study the performance of this inversion algorithm. The mean relative error measured by (20) between the ground truth and the reconstructed signal is $0.0298$. In particular, $90\%$ of the randomly picked window sets yield signal reconstructions with relative error less than $0.0320$. We also plot the reconstruction results from the uniformly chosen windows and one realization of the randomly selected windows in Figure 5.
@@ 4.2.2 Recovery of a multi-modal function with noisy measurements
Figure 7: Multi-modal function example: the noise-free measurement of $|V_{g}f|$ for $g=\varphi^{1}$ (left) and the noisy data used in phase retrieval (right).
We consider a multi-modal function plotted in Figure 6a, a linear combination of time-shifts of $\varphi^{4}$, the dilated Gauss function. The absolute value of its ambiguity function in log scale is shown in Figure 6b. This signal is known to cause stability issues in STFT phase retrieval [2, Sec. 6]. The measurement is contaminated by a multiplicative noise following the normal distribution $\mathcal{N}(1,0.05^{2})$; see Figure 7 for an example.
We consider the cutoff value $\varepsilon=0.5$ to define the essential support of the ambiguity function in Algorithms 1 and 2. Since $\varepsilon$ is large, we need to measure windows whose ambiguity functions have broader support in the time-frequency domain to perform the inversion. We first use fractionally Fourier-transformed versions of the dilated Gauss function $\varphi^{50}$ as the measuring windows. Since the essential support of these ambiguity functions is thin and long, we need to take more windows at different angles to achieve broad coverage of the joint essential support in the time-frequency domain: we use $80$ angles with $\alpha_{j}=(j-1)\pi/40$, $j=1,\ldots,80$. The recovery is shown in Figure 8, compared with the case using the standard Gauss as the measuring function. As seen in Figure 6b, the ambiguity function of the ground truth has almost disjoint support in the time-frequency domain. It thus cannot be stably recovered by a single Gauss window, as shown in Figure 8b, while the multiple fractional window strategy provides a stable recovery even under noisy measurement; see Figure 8a.
@@
Next, we also use Hermite functions as the measuring windows. In Figure 9a, we use $h_{50}$ and $h_{100}$, the Hermite functions of degree $50$ and $100$, to perform STFT phase retrieval following Algorithm 2. We recover most features of the multi-modal function with minimal impact from the measurement noise illustrated in Figure 7. We also compare the inversion result using the single window $h_{100}$ in Figure 9b. Although $\mathcal{A}h_{100}$ has extensive overall coverage in the time-frequency domain, it also has $100$ circular rings that are excluded since their absolute value is below the cutoff value $\varepsilon$. As a result, the reconstruction is affected by those missing data. The ambiguity function $\mathcal{A}h_{50}$ of the additional window $h_{50}$ makes up most of those “missing” rings, so the reconstruction is much better.
The large cutoff value $\varepsilon$ mitigates the impact of noise, while using multiple windows improves the coverage in the time-frequency domain of the joint essential support of the corresponding ambiguity functions. This is evident from both using multiple fractionally Fourier-transformed dilated Gauss functions and Hermite functions of various degrees.
@@
Figure 9: Multi-modal function example: (a) Recovery using two Hermite windows $h_{n}$ where $n=50,100$, and the relative error is $0.0899$; (b) Recovery using the single Hermite window $h_{100}$, and the relative error is $0.4403$ measured by (20). The true signal is plotted in blue.

### v2

@@ 3.1.3 Total error
$\quad+\left(\frac{1}{2(k\,k^{\eta})^{3/2}}\cdot\left\|A\right\|_{L^{p,1}(\mathbb{R}^{2})}+\frac{1}{k^{\eta}}\right)\cdot\frac{1}{\varepsilon}\max(C_{\mathbb{R},\Omega},C_{p})\left\|\sum_{j=1}^{N}\eta_{j}\right\|_{{L^{1}}(\mathbb{R}^{2})}$
(17) $\leq C\cdot\left(\left\|\mathcal{A}f\right\|_{L^{p,1}(\mathbb{R}^{2}\backslash\Omega)}+\left\|\mathcal{A}f(0,\cdot)\right\|_{L^{1}((\mathbb{R}^{2}\backslash\Omega)\cap(\left\{0\right\}\times\mathbb{R}))}+\frac{1}{\varepsilon}\left\|\sum_{j=1}^{N}\eta_{j}\right\|_{{L^{1}}(\mathbb{R}^{2})}\right).$
@@
Now, consider a fixed $\varepsilon$. If we only use a single window function, i.e., $N=1$, then the set $\Omega$ can be significantly smaller compared to the case with multiple windows; see Figs. 1 and 2 for illustrations. This again sheds lights on the fact that we achieve better stability by using more windows. In this way, we can significantly reduce the magnitude of the first two terms in (17) while keeping the last term constant.
@@ 3.2 Stability of Algorithm 2
In Algorithm 2, the reconstruction is done separately on each $\Omega_{j}$, using $F_{j}$, $F_{j}^{\eta}$ for computing $A$, $A^{\eta}$, respectively. Note that in Algorithm 2, the sets $\Omega_{j}$’s are constructed to be pairwise disjoint. Again, $A$ and $A^{\eta}$ will be set to zero outside of $\Omega$ and we obtain the following error bound for the data noise propagation:
@@
Let $A,A^{\eta}$ be the reconstructions obtained from $F_{j},$ $F_{j}^{\eta},$ respectively, $j=1,\dots,N$, and for some $\varepsilon>0$, let $\Omega_{j}$, pairwise disjoint, be chosen such that $\left|\mathcal{A}h_{n_{j}}(x,y)\right|>\varepsilon$ for all $(x,y)\in\Omega_{j}$ for all $j\in\{1,\dots,N\}$. Then,
$\left\|A-A^{\eta}\right\|_{L^{p,1}(\mathbb{R}^{2})}$$\leq$ $\frac{C_{p}}{\varepsilon}\max_{j=1,\dots,N}\left\|\eta_{j}\right\|_{{L^{1}}(\mathbb{R}^{2})},$
@@
$\left\|A-A^{\eta}\right\|_{L^{p,1}(\mathbb{R}^{2})}\leq\sum_{j=1}^{N}\left\|\mathds{1}_{\Omega_{j}}\frac{T\mathcal{F}(F_{j}-F_{j}^{\eta})}{\mathcal{A}h_{n_{j}}}\right\|_{L^{p,1}(\mathbb{R}^{2})}\leq\sum_{j=1}^{N}C_{j,p}\cdot\left\|\mathds{1}_{\Omega_{j}}\frac{T\mathcal{F}(F_{j}-F_{j}^{\eta})}{\mathcal{A}h_{n_{j}}}\right\|_{L^{\infty}(\mathbb{R}^{2})},$
where $C_{j,p}:=\|\mathds{1}_{\Omega_{j}}\|_{L^{p,1}(\mathbb{R}^{2})}<\infty$ and instead of (10), we now have the similar bound
$\left\|A-A^{\eta}\right\|_{L^{p,1}(\mathbb{R}^{2})}$ $\leq\sum_{j=1}^{N}\frac{C_{j,p}}{\inf_{\Omega_{j}}|\mathcal{A}h_{n_{j}}|}\cdot\left\|\mathcal{F}\eta_{j}\right\|_{L^{\infty}(\mathbb{R}^{2})}$
$\leq\frac{1}{\varepsilon}\sum_{j=1}^{N}C_{j,p}\cdot\left\|\eta_{j}\right\|_{{L^{1}}(\mathbb{R}^{2})}\leq\frac{C_{p}}{\varepsilon}\max_{j=1,\dots,N}\left\|\eta_{j}\right\|_{{L^{1}}(\mathbb{R}^{2})},$
where we have used $\sum_{j=1}^{N}C_{j,p}=C_{p}$ (since $\Omega_{1},\dots,\Omega_{N}$ are pairwise disjoint and $\cup_{j=1}^{N}\Omega_{j}=\Omega$), and the Hausdorff–Young inequality. Similar to (11), we can further derive
@@
With Lemma 3.9, the propagation of data noise can be estimated by
$\left\|\widetilde{f}-\widetilde{f\,}^{\eta}\right\|_{L^{p}(\mathbb{R})}$ $\leq\left(\frac{1}{2(k\,k^{\eta})^{3/2}}\cdot\left\|A\right\|_{L^{p,1}(\mathbb{R}^{2})}+\frac{1}{k^{\eta}}\right)\cdot\frac{1}{\varepsilon}\max(C_{\mathbb{R},\Omega},C_{p})\max_{j=1,\dots,N}\left\|\eta_{j}\right\|_{{L^{1}}(\mathbb{R}^{2})}.$
@@
$\quad+\left(\frac{1}{2(k\,k^{\eta})^{3/2}}\cdot\left\|A\right\|_{L^{p,1}(\mathbb{R}^{2})}+\frac{1}{k^{\eta}}\right)\cdot\frac{1}{\varepsilon}\max(C_{\mathbb{R},\Omega},C_{p})\max_{j=1,\dots,N}\left\|\eta_{j}\right\|_{{L^{1}}(\mathbb{R}^{2})}$
(18) $\leq C\cdot\left(\left\|\mathcal{A}f\right\|_{L^{p,1}(\mathbb{R}^{2}\backslash\Omega)}+\left\|\mathcal{A}f(0,\cdot)\right\|_{L^{1}((\mathbb{R}^{2}\backslash\Omega)\cap(\left\{0\right\}\times\mathbb{R}))}+\frac{1}{\varepsilon}\max_{j=1,\dots,N}\left\|\eta_{j}\right\|_{{L^{1}}(\mathbb{R}^{2})}\right),$
@@
The third term on the RHS is the only difference between (17) and (18). The one in (18) depends on the maximum noise power, while the one in (17) involves the sum of the noise for different window functions. Their differences result from the distinctions between Algorithms 1 and 2 regarding how the multi-window measurements are utilized in the STFT phase retrieval.
@@ 4.1 Discretization
(b) Recovery with a standard Gauss window, $\varepsilon=10^{-3}$
(c) Recovery with a standard Gauss window, $\varepsilon=10^{-6}$
(d) Recovery with a standard Gauss window, $\varepsilon=10^{-9}$
Figure 4: (a) Recovery using $40$ fractional Fourier transformed dilated Gauss windows at various angles; (b)-(d) Recovery using a single standard Gauss window with the cut-off $\varepsilon=10^{-3},10^{-6},10^{-9}$, respectively. The ground truth is plotted in blue.
@@
Figure 5: (a) Recovery using $6$ Hermite windows $h_{n}$, where $n=0,10,20,30,40,50$; (b) Recovery using $6$ Hermite windows $h_{n}$ with the index $n$ randomly selected between 00 and $50$, resulting in a set of indices $\{18,20,22,28,36,47\}$. The ground truth is shown in blue, the reconstructions obtained using Algorithm 2 are depicted as red dashed curves, and the absolute errors in the real components are illustrated with black solid lines.
@@ 4.2.1 Chirp signal recovery
As our first inversion example, we consider the chirp signal shown in Fig. 3a. As illustrated in Fig. 3b, its ambiguity function has extensive support in the time-frequency domain, making it challenging to reconstruct using a single window of the standard Gauss function whose ambiguity function has a very small essential support as seen in Fig. 1a. We set $\varepsilon=10^{-3}$ as a cut-off value, as used in Algorithms 1 and 2 to prevent the instability from dividing the phaseless measurements by a small value.
@@
We first use $40$ fractional Fourier transformed dilated Gaussian windows to perform the reconstruction following Algorithm 1, i.e., $\mathcal{F}_{\alpha_{j}}\varphi^{a}$ with $a=15$ and $\alpha_{j}=(j-1)\pi/40$, $j=1,\ldots,40$. The relative error based on formula (21) between the ground truth and the reconstructed signal is $0.0084$. We also use the standard Gauss function $\varphi^{1}$ as a single window function for comparison. The relative error is $0.7928$. After finding the global phase factor, we plot the reconstructions under different measuring window schemes in Fig. 4. More measuring windows yield a much more extensive joint coverage of the time-frequency domain, resulting in better signal reconstruction. The ambiguity function corresponding to the standard Gauss window has a very small essential support, as defined by the set $S_{n}^{\varepsilon}$ (see Fig. 1a), in contrast to the ambiguity function of the chirp signal, illustrated in Fig. 3b. Due to this limited coverage, reconstruction with the standard Gauss window only partially recovers the original signal for a cut-off value of $\varepsilon=10^{-3}$ (see Fig. 4b). As $\varepsilon$ is decreased to $10^{-6}$ and $10^{-9}$, more parts of the signal are recovered, as shown in Figs. 4c4d. However, the reconstruction remains incomplete even when $\varepsilon$ is the machine precision due to the inherently limited support of the ambiguity function associated with the standard Gauss window.
Next, we use multiple Hermite windows to perform the recovery. As demonstrated in Fig. 2, the essential support $S_{n}^{\varepsilon}$ of the ambiguity functions for Hermite functions of degree $n$ does not vary significantly between lower and moderate degrees, such as $n=1$ and $n=5$. However, as the degree $n$ increases, specific circular gaps may emerge in the set $S_{n}^{\varepsilon}$, particularly at higher degrees (e.g., $n=10$). These gaps motivate the need for additional windows to ensure sufficient coverage of the joint support. We propose selecting Hermite windows with degrees spaced at regular intervals, such as every 5 degrees (e.g., $n=0,5,10,\dots$). Intuitively, the degrees should be chosen to balance sufficient coverage of the joint support $\cup_{n\in I}S_{n}^{\varepsilon}$ and computational efficiency. A practical approach is to numerically evaluate the set $\cup_{n\in I}S_{n}^{\varepsilon}$ for the chosen degrees $I$ to verify that the joint support is appropriately large. Alternatively, one could randomly select degree indices $I$ within the range 00 (corresponding to the standard Gauss window) to a maximum degree $L$ and evaluate $\cup_{n\in I}S_{n}^{\varepsilon}$ to check for sufficient coverage. The random sampling method can sometimes reveal nontrivial degree combinations that improve reconstruction quality.
Consider $6$ Hermite function $h_{n}$ with $n\in\{0,10,20,30,40,50\}$. We then follow Algorithm 2 to perform the inversion. The misfit between the reconstructed signal and the ground truth measured by (21) is $0.0025$. On the other hand, we want to test the idea of randomly choosing window functions. We draw $6$ windows uniformly from the zeroth to the fiftieth degree Hermite functions, and use them to generate the phaseless data and perform the STFT phase retrieval following Algorithm 2. We repeat this procedure $100$ times to study the performance of this inversion algorithm. The mean relative error measured by (21) between the ground truth and the reconstructed signal is $0.0298$. In particular, $90\%$ of the randomly picked window sets yield signal reconstructions with relative error less than $0.0320$. We also plot the reconstruction results from the uniformly chosen windows and one realization of the randomly selected windows in Fig. 5.
@@ 4.2.2 Recovery of a multi-modal function with noisy measurements
Figure 7: Multi-modal function example: the noise-free measurement of $|V_{g}f|$ for $g=\varphi^{1}$ (left), the noisy data used in phase retrieval (middle) and the absolute error between the noise-free measurement and the noisy one (right).
We consider a multi-modal function plotted in Fig. 6a, a linear combination of time-shifts of $\varphi^{4}$, the dilated Gauss function. The absolute value of its ambiguity function in log scale is shown in Fig. 6b. This signal is known to cause stability issues in STFT phase retrieval [2, Sec. 6]. The measurement is contaminated by a multiplicative noise following the normal distribution $\mathcal{N}(1,0.05^{2})$; see Fig. 7 for an example where we plotted the noise-free (left) and noisy signal (middle), as well as the absolute error between the two (right).
We consider the cut-off value $\varepsilon=0.5$ to define the essential support of the ambiguity function in Algorithms 1 and 2. Since $\varepsilon$ is large, we need to measure windows whose ambiguity functions have broader support in the time-frequency domain to perform the inversion. We first use fractional Fourier transformed versions of the dilated Gauss function $\varphi^{50}$ as the measuring windows. Since the essential support of these ambiguity functions is thin and long, we need to take more windows at different angles to achieve broad coverage of the joint essential support in the time-frequency domain: we use $80$ angles with $\alpha_{j}=(j-1)\pi/40$, $j=1,\ldots,80$. The recovery is shown in Fig. 8, compared with the case using the standard Gauss as the measuring function. As seen in Fig. 6b, the ambiguity function of the ground truth has almost disjoint support in the time-frequency domain. It thus cannot be stably recovered by a single Gauss window, as shown in Fig. 8b, while the multiple fractional window strategy provides a stable recovery even from noisy measurements; see Fig. 8a. Similarly, the reason why Fig. 6b only recovers the ground truth over a small portion of the domain is that the ambiguity function associated with the standard Gauss window has a very limited essential support in the time-frequency domain compared to the ambiguity function of the ground truth signal (see Figs. 1a and 6b). While decreasing the cut-off value $\varepsilon$ may improve the reconstruction, it also reduces the stability of the algorithm when using the standard Gauss window.
@@
Next, we also use Hermite functions as the measuring windows. In Fig. 9a, we use $h_{50}$ and $h_{100}$, the Hermite functions of degree $50$ and $100$, to perform STFT phase retrieval following Algorithm 2. We recover most features of the multi-modal function with minimal impact from the measurement noise illustrated in Fig. 7. We also compare the inversion result using the single window $h_{100}$ in Fig. 9b. Although $\mathcal{A}h_{100}$ has extensive overall coverage in the time-frequency domain, it also has $100$ circular rings that are excluded since their absolute value is below the cutoff value $\varepsilon$. As a result, the reconstruction is affected by those missing data. The ambiguity function $\mathcal{A}h_{50}$ of the additional window $h_{50}$ makes up most of those “missing” rings, so the reconstruction is much better.
The large cut-off value $\varepsilon$ mitigates the impact of noise, while using multiple windows improves the coverage in the time-frequency domain of the joint essential support of the corresponding ambiguity functions. This is evident from both using multiple fractional Fourier transformed dilated Gauss functions and Hermite functions of various degrees.
@@
Figure 9: Multi-modal function example: (a) Recovery using two Hermite windows $h_{50}$ and $h_{100}$, where the relative error is $0.0881$; (b) Recovery using the single Hermite window $h_{100}$, and the relative error measured by (21) is $0.4417$. The ground truth is shown in blue, the reconstructions obtained using Algorithm 2 are depicted as red dashed curves, and the absolute errors in the real components are illustrated with black solid lines.